OpenStudy (anonymous):

I need to solve this problem in my exam: Suppose that the dollar cost of producing x radios is c(x) = 400 + 20x - 0.2x^2. Find the average cost per radio of producing the first 35 radios.

OpenStudy (anonymous):

A) $13.00 B) -$6.00 C) $455.00 D) $855.00

OpenStudy (asnaseer):

you need to work out the integral of c(x) from x=1 to 35 and then divide that answer by 35.

OpenStudy (anonymous):

c(x) = 400 + 20x - 0.2x^2. the averge cost is C= c(x)/x=400/x +20 -0.2x the ave cost of producing 35 radios is C(35)=400/35 +20 -0.2(35)= 24.4 for the first 35 radios ans.......

OpenStudy (anonymous):

24.43 (approx)